OpenStudy (anonymous):

Hanna finds 5 8 of 64. Which decimal is equivalent to 5 8 ? A 0.125 B 0.58 C 0.625 D 1.6

OpenStudy (mathmale):

I'd suggest you write "five eighths" as 5/8.\[5/8 = \frac{ 5 }{ 8 }=??\] Use a calculator to find the decimal equivalent of this fraction.

OpenStudy (anonymous):

? o.0 I am getting confuse

OpenStudy (anonymous):

\[as~\frac{ 5 }{ 8 } = 5 \div 8 = ? \]

OpenStudy (anonymous):

thanks both ;)

OpenStudy (mathmale):

Damara: Do you have a calculator? If so, please enter 5, then the division symbol, and then 8, and see what decimal fraction results. It's important that you know how to do this.